# PinRec: Outcome-Conditioned, Multi-Token Generative Retrieval for Industry-Scale Recommendation Systems

### v4

@@ 6.1.2. Metrics
We use engagement actions such as repins, grid clicks, and time spent. Additionally, we count “fulfilled sessions,” where the user takes at least one action in the session, and “unfulfilled sessions,” where the user does not take any action.
6.1.3. Results
In Figure 7, we visualize a Pareto front of the repin rate (relative to impressions) and grid click rate for the PinRec candidate generator by itself, with absolute numbers hidden due to non-public information. By reducing the grid click budget and correspondingly increasing the repin budget, we see the expected result as the grid click rate drops and repin rate increases. Similarly, we observe a significant correlation between the budgets and the overall A/B experiment metric lifts for Homefeed grid clicks and repins in Figure 8 in Appendix.
(a) PinRec-OC
(b) PinRec-{MT, OC}
Figure 7. Visualization of Pareto fronts for online rate of grid click versus repin for items recommended only by PinRec variants, with error bars denoted in grey.
For each PinRec variant, we selected the variant with the best fulfilled sessions and measured the percentage lift in metrics across Homefeed and the entire site (Table 2). Relative to the control group, the unconditioned variant of PinRec, PinRec-UC, demonstrates improvements in unfulfilled sessions and in Homefeed and site-wide grid clicks. In addition to those lifts, the outcome-conditioned variant, PinRec-OC, shows significant lift in site-wide fulfilled sessions. The multi-token variant, PinRec-{MT, OC}, significantly increased successful site sessions, time spent on site, and site-wide grid clicks.
@@ 6.1.3. Results
Type Metric UC OC {MT, OC}
Overall Fulfilled Sessions +0.02% +0.21% +0.28%
Metrics Time Spent -0.02% +0.16% +0.55%
Unfulfilled Sessions -0.28% -0.89% -0.78%
Action Site-wide Grid Clicks +0.58% +1.76% +1.73%
Metrics Homefeed Grid Clicks +1.87% +4.01% +3.33%
6.2. Search
On the Search surface, we evaluate the unconditioned variant of PinRec. As shown in Table 3, PinRec-UC improves search metrics significantly relative to a control group, with increased session-wide search fulfillment (sessions resulting in positive interactions like repins or long clicks). When deployed for promoted product retrieval, the same embeddings achieve a 1.83% reduction in CPA and 1.87% increase in shopping conversion volume in the US, indicating improved relevance and user engagement with sponsored content.
@@ 6.2. Search
Metric Type Metric PinRec-UC
Overall Metrics Search Fulfillment Rate +2.27%
Action Metrics Search Repins +4.21%
Search Grid Clicks +3.00%
6.3. Related Pins
On the Related Pins surface, we evaluate the outcome-conditioned variant of PinRec. As shown in Table 4, PinRec-OC improves Related Pins metrics significantly relative to a control group. We observe session-wide increase in fulfilled sessions and increase in sitewide engagement.
@@ 6.3. Related Pins
Overall Fulfilled Sessions +0.26%
Metrics Time Spent +0.30%
Unfulfilled Sessions -1.07%
Action Site-wide Grid Clicks +0.50%
Metrics Sitewide Repins +0.63%
@@ 7. Conclusion
We introduced PinRec, a generative retrieval system developed for Pinterest’s Homefeed, Search and Related Pins feed. To the best of our knowledge, this paper represents the first comprehensive study on deploying generative retrieval in a web-scale application. The incorporation of two key innovations — outcome-conditioned generation and temporal multi-token inference — enabled PinRec to enhance several business metrics without substantial cost increases.
Several promising future directions can be pursued. First, applying generative approaches to ranking, as explored by Zhai et al. (HSTU24), could be beneficial. Second, optimizing input sequences to the model, akin to prompt engineering in large language models, is another avenue worth exploring. Third, developing language model alignment techniques similar to Ouyang et al. (PPO22) could enhance the use of engagement signals from users. Overall, this paper highlights a promising avenue for utilizing generative retrieval in large-scale recommender systems.
@@ A.1.1. HSTU Architecture
Though PinRec relies primarily on the well-known transformer architecture, we note that our technique is completely architecture-agnostic. Specifically, outcome conditioning and multi-token prediction can be applied regardless of the specific architecture (e.g., transformer blocks) or algorithms used (e.g., attention), allowing them to be generic for any future use case of generative retrieval despite architectural or algorithmic improvements. To demonstrate this, we reproduce significant gains over baselines using the HSTU architecture (HSTU24), as opposed to the multi-head attention-based causal transformer. Note that we do not leverage the ranking task or any sequence sampling explored in (HSTU24) as it is not relevant for our specific use case. For simplicity, we reproduce results on the outcome conditioned variant of PinRec only, as we believe similar results can be extrapolated to other variants.
@@ A.1.2. Ablating Negative Sampling
Following from Pinnerformer22, we perform ablations on negative sampling, as it is a critical component of the chosen sampled softmax loss function. Though we ultimately end up with similar conclusions as Pinnerformer22, we show these findings for maximum clarity regarding our design choices.
@@ A.1.3. Ablating Parameter Scaling
One of the recent explorations within recommendation systems has been scaling to improve performance, in an attempt to replicate the success of these efforts for language modeling. To explore this hypothesis, we scale the number of parameters within PinRec across two fronts: the “vocabulary” and the model itself. While model scaling is clear and involves scaling up the embedding size, the number of transformer layers, and so on, scaling the vocabulary for an embedding-based generative retrieval technique involves improving the input representations. Specifically, as mentioned in the main text, we implement “ID” embeddings for pins, that effectively serve as additional representation for the “pin vocabulary”.
Model Parameters
@@
Though we observe some gains in Table 8 from scaling the number of transformer parameters to around 1 billion (XL-sized), these become intractable to serve in real time, requiring offline batch inference or significantly higher cost and more optimization. For the remaining sizes, we observe insignificant lifts (note that less than 1% in lift magnitude can be often explained through noise). Hence, we opt to not use larger sized transformers.
Vocabulary Parameters.
As presented in the main text, the ID embeddings consist of roughly 10 billion parameters. These are a core component of the model since they provide additional representational power for the most important entity type at Pinterest: Pins. To ablate the importance of these ID embeddings, we examine the performance in offline recall across all surfaces with and without ID embeddings. For this analysis, we focus on Homefeed (the most critical surface) and present results aggregated across all surfaces as well in Table 9.
Table 9. Percentage lift through adding ID embeddings (10 billion parameters) over no ID embedding-based vocabulary.
Surface % Lift in Recall @ 10
Homefeed +14.0%
All +14.4%
Evidently, the improvement in recall is substantial and indicates that the scaling of vocabulary to learn an embedding for each unique ID (roughly, as it is hash-based) rather than only content-based features is beneficial.
@@ A.1.4. Tail User and Item Analysis
In terms of new users, PinRec-UC/OC lift grid clicks and outbound clickthroughers (outbound CTs) in 10(a). Critically, the lift for outcome conditioning-based modeling is either neutral or positive with respect to both control and the unconditioned variant of PinRec, demonstrating that outcome conditioning does not suffer on such cohorts of users. In fact, across new users, OC performs better on outbound CTs.
@@
(b) PinRec-{MT, OC}
Figure 8. Visualization of Pareto fronts for online lift % in Homefeed grid clicks versus repins for PinRec variants.
A.1.5. Serving Efficiency
One of the critical aspects of our proposed generative retrieval technique is an effort to maintain serving efficiency. With our custom infrastructure built from off-the-shelf serving technologies such as NVIDIA Triton Inference Server and optimization techniques like torch.compile, we achieve a system that can be served in real-time. In this section, we delve deeper into the memory and time requirements, comparing our method with prior methods deployed in production.
On the memory front, all variants of PinRec (unconditioned, outcome-conditioned with or without multi-token generation) use about 1.6GiB of GPU memory including KVCache and CUDA graphs, without any significant memory usage increase for multi-token prediction or outcome conditioning. Excluding the KVCache and CUDA graphs, we use roughly the same amount of memory (2̃50 MiB) relative to other techniques (i.e., two-tower models).
In terms of latency, at 80 QPS, the online and offline latency for PinRec-OC/OC, MT are p50 40ms/p90 65ms (6 steps). On the other hand, across only 3 steps for PinRec-UC, we observe a latency of p50 23ms and p90 46ms (note that this produces significantly less embeddings and less compressed embeddings than either of the above techniques, though it is lower latency). Though the latency is roughly 3-4x higher (p50) than a traditional two-tower solution, we note that parallelizing it with other RPCs results in ¡1% added end-to-end increase in latency.
A.2. Visualization of Multi-Token Impact
(a) Recent user history manually specified to produce results in subfigures. In reverse chronological order. Note a mix of nail and tattoo-related queries and engagements.
(b) Results for PinRec-OC candidate generator for user history in subfigure above. Notably only includes Pins related to nails, the user’s most recent interaction.
(c) Results for PinRec-{MT, OC} candidate generator for user history in subfigure above. Includes a mix of Pins related to both tattoos and nails.
Figure 9. Comparison of results for a manually-specified user history.
(a) Recent user history manually specified to produce results in subfigures. In reverse chronological order. Note a mix of food and nursery-related queries and engagements.
(b) Results for PinRec-OC candidate generator for given user history. Notably, the model is attending more to recent food-related interactions than earlier nursery-related interactions.
(c) Results for PinRec-{MT, OC} candidate generator for given user history. Includes a mix of Pins related to nurseries and food.
Figure 10. Comparison of results for another manually-specified user history.
In this section, we present additional qualitative examples through visualizations of retrieved examples for sample user histories, comparing multi-token and non-multi-token variants. Through these visualizations, we are able to demonstrate a reasonable and consistent qualitative difference in the types of pins retrieved, justifying the improved session-wide and sitewide online metrics for PinRec-{MT, OC}.
Given a user history, we display two examples of retrieved pins in Figure 9 and Figure 10 for PinRec-{MT, OC} and PinRec-{OC}. Specifically, note that PinRec-{MT, OC} captures less-recent interactions for users compared to PinRec-OC. This is expected as the multi-token model predicts over a longer future time window and is less biased toward recency. PinRec-{MT, OC} can therefore produce more diverse results capturing more medium-term interests. On the other hand, since the PinRec-{OC} is naturally biased towards the short term, it fails to capture potential longer term trends that may induce engagement.
Note that these trends reflect the same takeaways from offline and online evaluation. Offline, we observe an improvement in diversity as the number of multi-token predictions is increased, which is consistently above the non-multi-token variant of PinRec. Online, we see that the “feed” presented by PinRec-{MT, OC} is of higher quality due to its diversity and visual appeal/aesthetic. These are critical qualities for a platform like Pinterest.

### v5

@@ 6.2.1. Metrics
We use engagement actions such as saves, grid clicks, and time spent. Additionally, we count “fulfilled sessions / fulfillment,” where the user takes at least one action in the session, and “unfulfilled sessions,” where the user does not take any action. Note that all reported metric in bold letters are statistically significant at the 5% level (two-sided test; $p<0.05$).
6.2.2. Home Feed
For Home Feed, we evaluated both PinRec-UC and PinRec-OC (Table 4). While PinRec-UC increases grid clicks, PinRec-OC delivers larger gains in downstream actions and in fulfilled sessions, demonstrating the benefit of outcome-conditioning.
@@ 6.2.2. Home Feed
Metric Type Metric UC OC
Overall Metrics Fulfilled Sessions +0.02% +0.21%
Action Site-wide Grid Clicks +0.58% +1.76%
Metrics Home Feed Grid Clicks +1.87% +4.01%
6.2.3. Related Pins
Table 5 shows the metric lift from the PinRec-OC on the Related Pins surface, with session-wide gains in fulfilled sessions and sitewide engagement. We also see cross-surface impact in Search: fulfillment increases by +0.42% and saves by +0.39%. PinRec did not change Search retrieval or ranking; these gains are driven indirectly by better Related Pins recommendations leading to more search actions, highlighting PinRec ’s cross-surface learning.
Metric Type Metric PinRec-OC
Overall Fulfilled Sessions +0.21%
Metrics Time Spent +0.38%
Action Site-wide Grid Clicks +0.76%
Metrics Sitewide Saves +0.26%
6.2.4. Search
Table 6 shows the results on Search A/B test. PinRec-OC improves search metrics significantly relative to a control group, with increased session-wide search fulfillment (sessions resulting in positive interactions like saves or long clicks).
@@ 6.2.4. Search
Metric Type Metric PinRec-OC
Overall Metrics Search Fulfillment Rate +2.24%
Action Metrics Search Saves +3.88%
Search Shares +5.30%
6.3. Zero-shot Retrieval
PinRec represents items and queries using shared, broadly available feature backbones (e.g., OmniSage (Badrinath et al., 2025) and OmniSearchSage (Agarwal et al., 2024)). Because these features are defined for a wide range of Pinterest entities, PinRec can be applied to items or queries even when they are absent from the PinRec training corpus. We therefore evaluate zero-shot retrieval, asking whether PinRec can retrieve entities that never appear during PinRec training.
Zero-shot capability is important in practice because Pinterest inventory and user intent evolve continuously: new Pins and ads are created every day, and new or tail queries emerge with limited interaction history. A retriever that generalizes zero-shot can serve such cold-start entities immediately, reducing reliance on frequent retraining and improving coverage for long-tail content.
After running PinRec inference for the PinRec candidate generator (CG), we take the model’s output embedding and perform nearest-neighbor search over other entity corpora:
  • Related Searches: retrieve search queries whose embeddings are closest to the user’s PinRec embedding.
  • Ads: retrieve ad Pins using the same nearest-neighbor retrieval approach.
For both the query and ads corpora, we precompute and index embeddings by running PinRec inference with their respective features described in Sec. 4.2.
Table 7 reports the results. Despite not being trained to predict engagement for these entities, PinRec retrieves highly relevant candidates (e.g., lower cost per ad action). Moreover, this retrieval is substantially cheaper in production because PinRec user-side inference is shared with organic CG. These results highlight the strong generalization of generative retrieval and support the case for a unified GR model.
Table 7. A/B experiments for PinRec as a zero-shot candidate generator.
@@ 6.3. Zero-shot Retrieval
Related Searches Fulfilled Sessions +1.66%
(Higher the better) Module Click Rate +4.04%
Ads Cost Per Acquisition -1.83%
(Lower the better) Cost Per Click -1.41%
@@ 7. Conclusion
We introduced PinRec, a unified generative retrieval system developed for Pinterest’s Homefeed, Search and Related Pins feed. To the best of our knowledge, this paper represents the first comprehensive study on deploying unified generative retrieval in a web-scale application. The incorporation of two key innovations — outcome-conditioned generation and surface-specific fine-tuning — enabled PinRec to enhance several business metrics across multiple applications.
Several promising future directions can be pursued. First, applying generative approaches to ranking, as explored by Zhai et al. (Zhai et al., 2024), could be beneficial. Second, optimizing input sequences to the model, akin to prompt engineering in large language models, is another avenue worth exploring. Third, developing language model alignment techniques similar to  He et al. (2025) could enhance the use of engagement signals from users. Overall, this paper highlights a promising avenue for developing unified generative retrieval for industrial recommender systems.
@@ A.1.1. HSTU Architecture
Though PinRec relies primarily on the well-known transformer architecture, we note that our technique is completely architecture-agnostic. Specifically, outcome conditioning can be applied regardless of the specific architecture (e.g., transformer blocks) or algorithms used (e.g., attention), allowing them to be generic for any future use case of generative retrieval despite architectural or algorithmic improvements. To demonstrate this, we reproduce significant gains over baselines using the HSTU architecture (Zhai et al., 2024), as opposed to the multi-head attention-based causal transformer. Note that we do not leverage the ranking task or any sequence sampling explored in (Zhai et al., 2024) as it is not relevant for our specific use case. For simplicity, we reproduce results on the outcome conditioned variant of PinRec only, as we believe similar results can be extrapolated to other variants.
@@ A.1.2. Ablating Negative Sampling
Following from Pancha et al. (2022), we perform ablations on negative sampling, as it is a critical component of the chosen sampled softmax loss function. Though we ultimately end up with similar conclusions as Pancha et al. (2022), we show these findings for maximum clarity regarding our design choices.
@@ A.1.3. Ablating Parameter Scaling
One of the recent explorations within recommendation systems has been scaling to improve performance, in an attempt to replicate the success of these efforts for language modeling. To explore this hypothesis, we scale the number of parameters within PinRec.
@@
Though we observe some gains in Table 11 from scaling the number of transformer parameters to around 1 billion (XL-sized), these become intractable to serve in real time, requiring offline batch inference or significantly higher cost and more optimization. For the remaining sizes, there was no improvement observed actually. Hence, we opt to not use larger sized transformers.
@@ A.1.4. Tail User and Item Analysis
In terms of new users, PinRec-UC/OC lift grid clicks and outbound clickthroughers (outbound CTs) in 12(a). Critically, the lift for outcome conditioning-based modeling is either neutral or positive with respect to both control and the unconditioned variant of PinRec, demonstrating that outcome conditioning does not suffer on such cohorts of users. In fact, across new users, OC performs better on outbound CTs.
@@
Figure 6. Visualization of Pareto fronts for online lift % in Home Feed grid clicks versus repins for PinRec
A.1.5. Embedding similarity threshold tuning
For the embedding similarity threshold for embedding compression in Sec. 4.4), we A/B tested values in the [0.75, 0.9] range for each surface. Our findings were: (1) results were qualitatively consistent across the different values, and (2) a threshold of 0.9 slightly outperformed other values for both Search and Related Pins, as reported in the paper. We will clarify this tuning procedure in the final version. We have not yet tuned this value for different user segments, but agree that this would be an interesting future research.
A.1.6. Temporal offset value tuning
Inference: We experimented with different values of $\delta$ for each action at inference time. For the results reported in this paper, we sampled repins and grid-clicks at $\delta\in\{0,25,50\}$ (i.e., 0, 25, and 50 discrete timesteps ahead). For all other outcomes, we sampled only the immediate next timestep ($\delta=0$). We also tried sampling every 5 timesteps from 0 to 50, but this generated too many embeddings for ANN search.
Training: We randomly selected 4 training targets with $\delta$ sampled from the interval $[0,25)$. For evaluation, we compared the model outputs at $\delta\in\{0,10\}$ and measured recall@10 for predicting the ground-truth action under different Timestep Gaps (TG). TG = 0 corresponds to immediate-term recall, while TG = 30 reflects long-term recall. Compared to $\delta=0$, $\delta=10$ achieved 60% higher long-term recall with only a 2% decrease in short-term recall, indicating a better overall trade-off; therefore, we use $\delta=10$.